# SAGE: Training Smart Any-Horizon Agents for Long Video Reasoning with Reinforcement Learning

### v1

@@ 4.3 Main Results
In Tab. 4, we compare our SAGE to Direct video reasoning methods, including models trained without RL post-training, like Qwen3-VL-4/8B-Instruct [qwen3-vl], and RL-tuned models, like Video-R1 [video-r1]. We also evaluate Agent systems like VideoMind [liu2025videomind] and VideoExplorer [videoexplorer].
Effective System Design. We separately evaluate the performance of our system with two API-based models SAGE-MM: Gemini-2.5-Flash [gemini25pushingfrontier] and GPT-4o [openai2024gpt4ocard]. For this setting, we use Gemini-2.5-Flash as the backend model for the ground-event and analyze tools; therefore, we denote the system as SAGE-Flash. We observe improvements of up to 3.2% over the base API models, validating the effectiveness of our system design.
@@
Effective Training Recipe. As shown in Tab. 4, our SAGE with a trained SAGE-MM achieves notable improvements across different base MLLMs. Specifically, SAGE surpasses Qwen2.5-VL-7B-Instruct by 4.8% overall, with substantial gains of +6.1% on open-ended and +7.0% on visual questions, underscoring the effectiveness of our training strategy. Interestingly, models such as Video-R1 [video-r1], VideoRFT [VideoRFT], and VideoExplorer [videoexplorer], despite employing finetuned Qwen2.5-VL-7B-Instruct backbones, underperform relative to the base model, particularly on open-ended questions. Moreover, as shown in the last row of Tab. 4, SAGE-Flash further improves upon SAGE by 3.8%, even outperforming the Gemini-2.5-Flash variant of SAGE-MM. This indicates that our finetuned SAGE-MM not only learns to invoke tools effectively but also benefits from more accurate tool outputs.
Additionally, we report results with Qwen2.5-VL-7B-Instruct based SAGE-MM on MINERVA [minerva], a complex video reasoning benchmark that covers domains such as sports, short films, and cooking videos. As shown in Tab. 5, our SAGE shows an improvement of 2.6% on long videos (duration >>600 seconds) compared to the base model while outperforming other reasoning models, validating the effectiveness of our approach for long video reasoning.
@@ 4.4 Ablations
VideoRFT-7B [VideoRFT] Direct 128 55.3 7.2
Video-R1-7B [video-r1] Direct 128 57.6 7.3
VideoMind-7B [liu2025videomind] Agent 50.0 24.7
LVAagent [chen2025lvagent] Agent 49.7 92.9
VideoChat-R1.5-7B [yan2025videochatr15] Agent 54.8 132.1
VideoExplorer-7B [videoexplorer] Agent 50.1 137.7
VideoAgent [fan2025videoagent] Agent 42.0 1445.0
@@
In Tab. 10, we compare the accuracy score and inference runtime per sample for our SAGE to other existing Direct [video-r1, VideoRFT] and Agent [videoexplorer, fan2025videoagent, chen2025lvagent] baselines and various frame-input setups of the baseline Qwen3-VL-8B-Instruct [qwen3-vl]. We observe that although the runtime of our SAGE is comparable to using 512 frames as inputs to Qwen3-VL-8B-Instruct, it shows far superior performance, while only being slower by about 1 second compared to other thinking Direct baselines. Moreover, our system is almost 3 times quicker than VideoMind [liu2025videomind], the quickest Agent baseline, demonstrating the superiority of our system design and training recipe for practical applications over existing systems.
The lower runtime of our framework compared to the Agent baselines is primarily due to the baselines’ system design, which involves heavy video preprocessing and excessive recurrent model calls. Notably, VideoAgent [fan2025videoagent] is slowed by a mandatory preprocessing phase in which every 2-second subclip undergoes multi-model analysis for metadata extraction, making it super slow for long videos. Similarly, VideoExplorer [videoexplorer] suffers from both an initial  30-second preprocessing delay, arising from dividing the video into multiple subclips for embedding computation, and an inference process involving multiple retrieval steps. Finally, VideoMind [liu2025videomind] inherently requires more model invocations. This increase can be traced to the system design, which requires repetitive invocations of the verifier module. The verifier is executed multiple times, once for each of the top five potential segments generated by the preceding grounder module, which slows the system.
@@ 5 Conclusion
Future Work. Looking ahead, training on data from broader domains to handle more use cases is a natural advancement. In addition, integrating more advanced agent-centric policy optimization algorithms [feng2025group, dong2025aepo, gspo] for RL presents a promising avenue. Finally, empowering the system to select the appropriate tools and synthesize new ones when necessary [prabhu2025waltwebagentslearn, vadar] represents an exciting direction.
@@ 6 Additional Ablations
In Tab. 12, we analyze the effect of eval mode with the trained Qwen3-VL-8B-Instruct SAGE-MM during inference. We find that the Agent mode performs better than Direct mode. Surprisingly, the Qwen3-VL [qwen3-vl] based model shows much better performance than the Molmo2 [molmo2] one with Direct eval mode which could be attributed to the two model families’ different abilities to learn information directly since all the models are trained under the Agent paradigm.
@@
Our SAGE is designed so that any MLLM with function-calling capabilities can be used as the SAGE-MM. In Tab. 13, we evaluate the base Qwen3-VL [qwen3-vl] models as SAGE-MM, without any finetuning. We observe that Qwen3-VL-4B-Instruct is not an effective orchestrator: it rarely engages in multi-turn reasoning and attains low accuracy, indicating that SFT is essential before applying RL.
Interestingly, the base Qwen3-VL-8B-Instruct model behaves differently. It is a noticeably stronger function caller, demonstrating a more reasonable balance between single-turn and multi-turn reasoning. This motivates us to apply RL directly on top of the base model to assess the importance of SFT for the 8B variant. Surprisingly, RL without SFT fails, i.e., the model collapses to single-turn reasoning. We hypothesize that this is due to the base model’s training objective, which strongly biases it toward directly producing final answers, making SFT necessary to incentivize [wang2025octothinker] any-horizon reasoning during RL. While it is possible that a heavily engineered RL recipe could overcome this, we do not pursue this direction, as SFT is far simpler and cheaper than extensive hyperparameter tuning during RL.
@@ 7 SAGE-Bench
As illustrated in Fig. 5, existing popular benchmarks [zhou2025mlvubenchmarkingmultitasklong, cheng2025video, hu2025videommmuevaluatingknowledgeacquisition, wu2024longvideobenchbenchmarklongcontextinterleaved, fu2025videommefirstevercomprehensiveevaluation] rely purely on multiple-choice questions (MCQs). In contrast, SAGE-Bench utilizes open-ended questions with an unbounded answer space, aligning more closely with practical user situations.
@@
While AI systems are ultimately intended for real-world deployment, existing benchmarks often include diagnostic questions to gauge models’ visual understanding, such as temporal ordering tasks in MLVU [zhou2025mlvubenchmarkingmultitasklong]. SAGE-Bench incorporates both diagnostic questions to test fundamental model capabilities and practical questions that users may have while watching entertainment videos, ensuring that our benchmark evaluates not only technical proficiency but also the model’s utility in practical scenarios.

### v2

@@ 4.3 Main Results
In Tab. 4, we compare our SAGE to Direct video reasoning methods, including models trained without RL post-training, like Qwen3-VL-4/8B-Instruct [42], and RL-tuned models, like Video-R1 [18]. We also evaluate Agent systems like VideoMind [30] and VideoExplorer [57].
Effective System Design. We separately evaluate the performance of our system with two API-based models SAGE-MM: Gemini-2.5-Flash [40] and GPT-4o [33]. For this setting, we use Gemini-2.5-Flash as the backend model for the ground-event and analyze tools; therefore, we denote the system as SAGE-Flash. We observe improvements of up to 3.2% over the base API models, validating the effectiveness of our system design.
@@
Effective Training Recipe. As shown in Tab. 4, our SAGE with a trained SAGE-MM achieves notable improvements across different base MLLMs. Specifically, SAGE surpasses Qwen2.5-VL-7B-Instruct by 4.8% overall, with substantial gains of +6.1% on open-ended and +7.0% on visual questions, underscoring the effectiveness of our training strategy. Interestingly, models such as Video-R1 [18], VideoRFT [46], and VideoExplorer [57], despite employing finetuned Qwen2.5-VL-7B-Instruct backbones, underperform relative to the base model, particularly on open-ended questions. Moreover, as shown in the last row of Tab. 4, SAGE-Flash further improves upon SAGE by 3.8%, even outperforming the Gemini-2.5-Flash variant of SAGE-MM. This indicates that our finetuned SAGE-MM not only learns to invoke tools effectively but also benefits from more accurate tool outputs.
Additionally, we report results with Qwen2.5-VL-7B-Instruct based SAGE-MM on MINERVA [32], a complex video reasoning benchmark that covers domains such as sports, short films, and cooking videos. As shown in Tab. 5, our SAGE shows an improvement of 2.6% on long videos (duration >>600 seconds) compared to the base model while outperforming other reasoning models, validating the effectiveness of our approach for long video reasoning.
Generalization to Other Benchmarks. We also evaluate on Video-MMMU [22] and Video-MME [20] in Tab. 6. Our SAGE-Flash outperforms baselines on Video-MMMU, demonstrating generalization to knowledge acquisition from videos. On Video-MME, the perception-centric nature of the benchmark means ground-event/extract-video-parts tools can hurt performance; disabling them recovers competitive results.
@@ 4.4 Ablations
VideoRFT-7B [46] Direct 128 55.3 7.2
Video-R1-7B [18] Direct 128 57.6 7.3
VideoMind-7B [30] Agent 50.0 24.7
LVAagent [3] Agent 49.7 92.9
VideoChat-R1.5-7B [53] Agent 54.8 132.1
VideoExplorer-7B [57] Agent 50.1 137.7
VideoAgent [17] Agent 42.0 1445.0
@@
In Tab. 11, we compare the accuracy score and inference runtime per sample for our SAGE to other existing Direct [18, 46] and Agent [57, 17, 3] baselines and various frame-input setups of the baseline Qwen3-VL-8B-Instruct [42]. We observe that although the runtime of our SAGE is comparable to using 512 frames as inputs to Qwen3-VL-8B-Instruct, it shows far superior performance, while only being slower by about 1 second compared to other thinking Direct baselines. Moreover, our system is almost 3 times quicker than VideoMind [30], the quickest Agent baseline, demonstrating the superiority of our system design and training recipe for practical applications over existing systems.
The lower runtime of our framework compared to the Agent baselines is primarily due to the baselines’ system design, which involves heavy video preprocessing and excessive recurrent model calls. Notably, VideoAgent [17] is slowed by a mandatory preprocessing phase in which every 2-second subclip undergoes multi-model analysis for metadata extraction, making it super slow for long videos. Similarly, VideoExplorer [57] suffers from both an initial  30-second preprocessing delay, arising from dividing the video into multiple subclips for embedding computation, and an inference process involving multiple retrieval steps. Finally, VideoMind [30] inherently requires more model invocations. This increase can be traced to the system design, which requires repetitive invocations of the verifier module. The verifier is executed multiple times, once for each of the top five potential segments generated by the preceding grounder module, which slows the system.
@@ 5 Conclusion
Future Work. Looking ahead, training on data from broader domains to handle more use cases is a natural advancement. In addition, integrating more advanced agent-centric policy optimization algorithms [19, 15, 60] for RL presents a promising avenue. Finally, empowering the system to select the appropriate tools and synthesize new ones when necessary [34, 31] represents an exciting direction.
@@ 6 Additional Ablations
In Tab. 13, we analyze the effect of eval mode with the trained Qwen3-VL-8B-Instruct SAGE-MM during inference. We find that the Agent mode performs better than Direct mode. Surprisingly, the Qwen3-VL [42] based model shows much better performance than the Molmo2 [10] one with Direct eval mode which could be attributed to the two model families’ different abilities to learn information directly since all the models are trained under the Agent paradigm.
@@
Our SAGE is designed so that any MLLM with function-calling capabilities can be used as the SAGE-MM. In Tab. 14, we evaluate the base Qwen3-VL [42] models as SAGE-MM, without any finetuning. We observe that Qwen3-VL-4B-Instruct is not an effective orchestrator: it rarely engages in multi-turn reasoning and attains low accuracy, indicating that SFT is essential before applying RL.
Interestingly, the base Qwen3-VL-8B-Instruct model behaves differently. It is a noticeably stronger function caller, demonstrating a more reasonable balance between single-turn and multi-turn reasoning. This motivates us to apply RL directly on top of the base model to assess the importance of SFT for the 8B variant. Surprisingly, RL without SFT fails, i.e., the model collapses to single-turn reasoning. We hypothesize that this is due to the base model’s training objective, which strongly biases it toward directly producing final answers, making SFT necessary to incentivize [49] any-horizon reasoning during RL. While it is possible that a heavily engineered RL recipe could overcome this, we do not pursue this direction, as SFT is far simpler and cheaper than extensive hyperparameter tuning during RL.
@@
Per-Tool Accuracy.
We report per-tool accuracy in Tab. 18, evaluating performance when only a single tool is available. We observe that extract-video-parts/ground-event perform the worst, attributed to their dependence on tools like analyze/extract-video-parts for local segment processing. The best individual tool performance comes from transcribe-speech, highlighting the importance of speech information for long video understanding.
all-tools 71.8 transcribe-speech 61.1 web-search/parse-website 58.6
analyze 58.6 extract-video-parts 50.2 ground-event 50.3 no-tools 53.0
Table 18: Per-tool performance comparison on SAGE-Bench.
Qualitative Failure Analysis.
We observe a few failure patterns of SAGE. First, SAGE may overcall tools when a tool invocation fails, resulting in retries that consume reasoning steps. Second, inaccurate temporal grounding can result in missing video segments, leading to incomplete context. Lastly, misleading transcript or web information can result in incorrect answers.
@@ 7 SAGE-Bench
As illustrated in Fig. 5, existing popular benchmarks [62, 9, 22, 51, 20] rely purely on multiple-choice questions (MCQs). In contrast, SAGE-Bench utilizes open-ended questions with an unbounded answer space, aligning more closely with practical user situations.
@@
While AI systems are ultimately intended for real-world deployment, existing benchmarks often include diagnostic questions to gauge models’ visual understanding, such as temporal ordering tasks in MLVU [62]. SAGE-Bench incorporates both diagnostic questions to test fundamental model capabilities and practical questions that users may have while watching entertainment videos, ensuring that our benchmark evaluates not only technical proficiency but also the model’s utility in practical scenarios.